How many mph am I going if I go 3/8 of a mile in 19 seconds?
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

360/19=18.95
3/8*18.95=7.1 mph

Answer by Alan3354(69443)

How many mph am I going if I go 3/8 of a mile in 19 seconds?
---------
= (3/8)/19 mi/sec*60 sec/min * 60 min/hr